If it takes 50 mL of 0.5 M KOH solution to completely neutralize 125 mL of nitric acid solution (HNO3), what is the concentration of the nitric acid solution?


1
Expert's answer
2022-01-24T23:12:02-0500

2KOH(aq) + H2SO4(aq) → K2SO4(aq) + 2H2O(l)


2mol KOH react with 1 mol H2SO4


Mol KOH in 50mL of 0.50M solution = 50mL / 1000mL/L × 0.50mol/L = 0.025 mol


This will react with 0.025/2 = 0.0125 mol H2SO4


Therefore 125mL of H2SO4 solution contains 0.0125 mol H2SO4

1.0L = 1000mL solution contains:

mol H2SO4 = 1000mL/L / 125mL * 0.0125mol = 0.1 mol

Molarity of H2SO4 solution = 0.1M